# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v5

@@ IV Existing QRNG Models based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging PhysRevLett.72.210 ; kim1999single . However, different experimental approaches PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval argillander2023quantum . In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
@@
If $\lambda$ is the wavelength of the photon wave and $n$ photons are present within $T$ time, the corresponding photon power would be
$P=\frac{n\hbar c}{\lambda T}=\frac{\mu\hbar c}{\lambda},$ (8)
where $\hbar$ is the Planck’s constant and $c$ is the speed of light in vacuum. Let $P_{a}$ be the photon power coming out from the attenuators. Then, the expected number of photons per unit time is given by
$\mu=\frac{P_{a}\lambda}{\hbar c}.$ (9)
This attenuated pulse is detected by the SPD. The time of this detection is registered to produce photon arrival time-based random numbers.
@@
Figure 3: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 00 if $\Delta t_{2i-1}<\Delta t_{2i}$ and $1$ if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins of equal length. If a photon has been detected in the $i$-th bin then $i-1$ will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then $i-1$ will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Fig. 3. In Fig. 3(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate the $i$-th bit of the random numbers as follows.
[…]
@@
But in this process 10.1063/1.2961000 , the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature 10.1063/1.2720728 ; WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
$P(t_{w}>t)$ $=P(\text{no photon in }(0,t])$
$=P(0,t)=e^{-\mu t}.$ (11)
Therefore, the random number generated with this procedure follows exponential distribution WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . Another QRNG scheme is proposed in 10.1063/1.2720728 , as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
Let $n$ photons be present within $T$ time after the attenuation. Since the photon count follows Poisson distribution (4), the probability, that all the $n$ photons are detected in the $i$-th interval, is given by
$P(i;all|n)$ $=\frac{P(n,|\tau_{i}|)P(0,\frac{i-1}{N}T)P(0,T-\frac{i}{N}T)}{P(n,T)}$
$=\frac{\frac{e^{-\mu|\tau_{i}|}(\mu|\tau_{i}|)^{n}}{n!}e^{-\mu\frac{i-1}{N}T}e^{-\mu(1-\frac{i}{N})T}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\frac{1}{N^{n}}.$ (12)
Here $|\tau_{i}|$ is the length of $i$-th time-bin, which equals $\frac{T}{N}$. Therefore, if only one photon is present within $T$ time after the attenuation then that photon would be detected in the $i$-th bin with probability $\frac{1}{N}$. Thus the photon arrival time follows random distribution over the time-bins. However, this is the case, when exactly one photon is present in $T$ time, and also when the detection is perfect. Since this is not a practical case, next we discuss the photon statistics for some practical scenarios.
V Our Proposal of Two-fold Statistical Method for Testing Quantumness
@@
In Section II we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance $=2$. But, it is not Poisson.
Also, in experimental methods the means and variances are calculated from finite data samples. The fineness of the sample cannot perfectly reflect the mean and variances of the underlying distribution. Therefore, the comparison between the mean and the variance with equality may not reflect the underlying distribution undoubtedly.
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. Numerous photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and experimentally demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years, providing viable tools for carrying out these tests. The statistical tools have been used in this section is discussed in C.
V.1 First Phase: Interval Estimation
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distribution can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sub-Poissonian or not by estimating the confidence interval for the mean of the underlying distribution as follows.
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$ hogg2019introduction , where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
@@ V.1 First Phase: Interval Estimation
Let $1-\epsilon$ be the confidence level. Suppose $y_{\epsilon}$ be such that
[…]
@@
[…]
distribution with confidence level $1-\epsilon$.
@@
V.2 Hypothesis Testing as an Alternative to Interval Estimation
To determine the quantumness one has to perform two consecutive hypothesis testing. For the first test,
@@ V.2 Hypothesis Testing as an Alternative to Interval Estimation
where $m$ is the mean of the actual distribution and $s^{2}$ is given by Eq. (14). For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test,
case 1: if $z\geq y_{\epsilon}$, where $y_{\epsilon}$ is given by Eq. (15), $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian.
case 2: if $z<y_{\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform second hypothesis testing with the same statistic. After the second test, if $z\geq y_{\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian.
Note that, this hypothesis testing is equivalent to the above interval estimation technique, and both results in the same decision.
V.3 Second Phase: Goodness-of-fit Test
@@
where $O_{n}$ denotes the frequency of $n$ number of photon in the data, $E_{n}$ denotes the frequency given by Poisson distribution (4). This test statistic follows $\chi^{2}$ distribution with $k=N-1$ degrees of freedom if all the parameters are known, and its density function is given by,
@@
[…]
Figure 4: Probability density function (22) for $\chi^{2}$ distribution. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by stars, and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
Figure 4 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. The area of the critical region is given by
@@
If the typical values of the parameters $\mu,T$ in Eq. (4) are known, we have $k=N-1$. However, if they are unknown, in that case, $k=N-2$. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if $p$-value is less than $\alpha$.
As in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test to ensure the completeness of the discussion.

### v6

@@ II Existing QRNG Models based on Coherent Pulse
Although a perfect single photon source ensures the quantum behaviour of light, it is well known that producing single photon states is experimentally challenging PhysRevLett.72.210 ; kim1999single . However, different experimental approaches PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval argillander2023quantum . In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 1.
@@
Several methods exist for generating random numbers from registered photon arrival times, illustrated in Fig. 2. In Fig. 2(I), given arrival times $\{t_{0},t_{1},\dots,t_{2i}\}$, two consecutive non-overlapping time intervals $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2}$ and $\Delta t_{2i}=t_{2i}-t_{2i-1}$ are used to generate the $i$-th bit as:
$\begin{cases}0,&\text{if }\Delta t_{2i-1}<\Delta t_{2i},\\ 1,&\text{if }\Delta t_{2i-1}>\Delta t_{2i}.\end{cases}$ (1)
This method 10.1063/1.2961000 ; Schranz2024 has low efficiency, producing one bit per two photon detections.
@@
Figure 2: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 0 if $\Delta t_{2i-1}<\Delta t_{2i}$ and 11 if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins of equal length. If a photon has been detected in the $i$-th bin then i1i-1 will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then i1i-1 will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
@@
Faster methods have been proposed 10.1063/1.2720728 ; WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 ; solymos2024correlation . In Fig. 2(II), time is divided into equal-length bins. If a photon is detected in the $i$-th bin, the value i1i-1 is taken as the random number. A new cycle begins after a fixed detector sleep time. Since photon detection follows a Poisson process, the probability that the waiting time $t_{w}$ between two photons exceeds $t$ is $P(t_{w}>t)=P(\text{no photon in }(0,t])=P(0,t)=e^{-\mu t}$, indicating an exponential distribution of outcomes WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 .
A third approach 10.1063/1.2720728 , shown in Fig. 2(III), uses an external clock divided into cycles of length $T$, smaller than the detector’s dead time. Each cycle is further divided into $N$ equal bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. If the first detected photon in a cycle arrives in bin $\tau_{i}$, the output is i1i-1. The detector remains inactive for the rest of the cycle.
@@
Let $n$ photons arrive within time $T$ after attenuation. Since the photon count follows Poisson distribution (47), the probability that all $n$ photons fall in bin $\tau_{i}$ is:
$P(i;\text{all}|n)$ $=\frac{P(n,|\tau_{i}|)\,P(0,\frac{i-1}{N}T)\,P(0,T-\frac{i}{N}T)}{P(n,T)}$
$=\frac{e^{-\mu|\tau_{i}|}(\mu|\tau_{i}|)^{n}/n!\cdot e^{-\mu\frac{i-1}{N}T}\cdot e^{-\mu(1-\frac{i}{N})T}}{e^{-\mu T}(\mu T)^{n}/n!}=\frac{1}{N^{n}},$ (2)
where $|\tau_{i}|=\frac{T}{N}$. Thus, if exactly one photon is present within $T$ and detection is ideal, it falls in any bin with equal probability $1/N$, producing uniformly distributed outcomes.
However, this ideal case assumes perfect single-photon input and detection, conditions rarely met in practice. In the following, we analyze photon statistics under more realistic experimental scenarios.
III Our Proposal of Two-fold Statistical Method for Testing Quantumness
@@
For the Poissonian distribution, the mean photon count is the same as its variance. However, it is interesting to note that mean == variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance =2=2. But it is not Poisson.
Also, in experimental methods, the means and variances are calculated from finite data samples. The fineness of the sample cannot perfectly reflect the mean and variance of the underlying distribution. In such a scenario, direct comparison between the mean and the variance may not always work. Suppose we are given a device to test whether the photons emitted from it are in the quantum regime. There are two possibilities – either the device is a single photon source with sub-Poissonian distribution or it is a source of coherent pulses with Poissonian distribution111very rarely it will be super-Poissonian. In the former case, the old method of point estimation will successfully predict the correct nature of the device. However, in the latter case, it will fail most of the time. This may lead to a wrong conclusion about the underlying distribution. For example, in the case of a device based on coherent pulse, which follows a Poisson distribution, a small error in the sample mean or in the sample variance may lead to a conclusion that the underlying distribution is sub-Poissonian, indicating a proper quantum source of light.
This motivates us to propose a two-fold method to test the quantumness of the photon source, as an answer to question Q1, whether the source is a purely single-photon source or a source of coherent pulses. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. The statistical tools used in this section is discussed in C.
III.1 First Phase: Interval Estimation
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distributions can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sub-Poissonian or not by estimating the mean and the variance of the underlying distribution. However, estimation of the values, known as point estimation, is not always exact ross2009introduction ; mood1974 . Thus, estimating two parameters, the mean and the variance, would increase the error. Therefore, we suggest estimation of an interval for one parameter with some confidence, called significance level. Then we can compare the other parameter with this interval.
First, we compare the variance with the interval of the mean. Here, we find an interval where the distribution mean would lie up to some given confidence and a value estimating the variance of the distribution. Then, we compare the estimated variance with the interval to decide the nature of the distribution. Suppose the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$ hogg2019introduction , where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
@@ III.1 First Phase: Interval Estimation
Let $1-\epsilon$ be the confidence level. Suppose $y_{\epsilon}$ is such that
$P(-y_{\epsilon}<Y<y_{\epsilon})\approx\frac{1}{\sqrt{2\pi}}\int_{-y_{\epsilon}}^{y_{\epsilon}}e^{-\frac{y^{2}}{2}}dy=1-\epsilon.$ (5)
@@
$\frac{\langle x\rangle}{s^{2}}\geq 1+\frac{y_{\epsilon}}{s\sqrt{n}}\implies$ $\text{sub-Poissonian distribution},$ (8)
$\frac{\langle x\rangle}{s^{2}}\leq 1-\frac{y_{\epsilon}}{s\sqrt{n}}\implies$ $\text{super-Poissonian distribution},$
with confidence level $1-\epsilon$.
The performance of this process depends on how the sample mean and sample variance converge to the distribution mean and the distribution variance as $n$ increases. The variance of the sample mean and the sample variance are given by $var[\langle X\rangle]=\frac{\sigma^{2}}{n}$ and $var[S^{2}]=\frac{1}{n}\left(\mu_{4}-\frac{n-3}{n-1}\sigma^{4}\right)$, respectively, where $\mu_{4}$ is the fourth order central moment defined as $\mu_{4}:=\langle(X-\langle X\rangle)^{4}\rangle$ ross2009introduction ; casella2002statistical ; mood1974 . As $\frac{var[\langle X\rangle]}{var[S^{2}]}\approx\mathcal{O}(\sigma^{2})$, relative convergence of the mean and the variance depends on the distribution variance. Thus, if $\sigma^{2}\gg 1$, this process may fail to identify the nature of the distribution.
To overcome the above limitation, we suggest an alternative approach to compare the mean with the interval of the variance. Here, we find an interval for the distribution variance up to a given confidence and compare the mean with it. Suppose the sample contains $n$ data points for the number of photon counts within a fixed time. Let $\langle x\rangle$ and $S^{2}$ be the sample mean and sample variance, respectively. Then $\langle x\rangle$ is an unbiased estimate for the distribution mean ross2009introduction ; casella2002statistical . It is also consistent if the variance is known. I $\sigma$ is the variance of the distribution, the probability distribution of the variance is given by $\frac{nS^{2}}{\sigma^{2}}$, which follows $\chi^{2}$ distribution. Then for the confidence level $1-\epsilon$, we have two numbers $\chi^{2}_{n-1,1-\epsilon/2}$ and $\chi^{2}_{n-1,\epsilon/2}$ such that
$P(\chi^{2}_{n-1,1-\epsilon/2}<\chi^{2}<\chi^{2}_{n-1,\epsilon/2})=1-\epsilon$ (9)
$\iff$ $P(\chi^{2}_{n-1,1-\epsilon/2}<\frac{nS^{2}}{\sigma^{2}}<\chi^{2}_{n-1,\epsilon/2})=1-\epsilon$
$\iff$ $P\left(\frac{nS^{2}}{\chi^{2}_{n-1,\epsilon/2}}<\sigma^{2}<\frac{nS^{2}}{\chi^{2}_{n-1,1-\epsilon/2}}\right)=1-\epsilon.$ (10)
Therefore, $\left(\frac{nS^{2}}{\chi^{2}_{n-1,\epsilon/2}},\frac{nS^{2}}{\chi^{2}_{n-1,1-\epsilon/2}}\right)$ is the confidence interval for the distribution variance up to the confidence interval $1-\epsilon$. Thus, for some given $\epsilon$, we can find $\chi^{2}_{n-1,1-\epsilon/2},\chi^{2}_{n-1,\epsilon/2}$ by Eq. (9) and conclude that
$\langle x\rangle>\frac{nS^{2}}{\chi^{2}_{n-1,1-\epsilon/2}}\implies$ $\text{sub-Poissonian distribution},$ (11)
$\langle x\rangle<\frac{nS^{2}}{\chi^{2}_{n-1,\epsilon/2}}\implies$ $\text{super-Poissonian distribution},$
with confidence level $1-\epsilon$.
@@
III.2 Hypothesis Testing as an Alternative to Interval Estimation
To determine the quantumness, one has to perform two consecutive hypothesis tests. For the first test,
@@ III.2 Hypothesis Testing as an Alternative to Interval Estimation
where $m$ is the mean of the actual distribution and $s^{2}$ is given by Eq. (4). For both tests, the test statistic would be the dispersion index Fisher1992 ; Alves2022 $\chi^{2}=\frac{(n-1)s^{2}}{\langle x\rangle}$. For the first test,
case 1: if $\chi^{2}\geq\chi^{2}_{n-1,1-\epsilon}$, where $\chi^{2}_{n-1,1-\epsilon}$ is the critical Chi-square value with n1n-1 degrees of freedom at confidence level $1-\epsilon$, $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian.
case 2: if $\chi^{2}<\chi^{2}_{n-1,1-\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform a second hypothesis test with the same statistic. After the second test, if $\chi^{2}\leq\chi^{2}_{n-1,\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian.
III.3 Second Phase: Goodness-of-fit Test
@@
where $O_{n}$ denotes the frequency of $n$ number of photon in the data, $E_{n}$ denotes the frequency given by Poisson distribution (47). This test statistic follows $\chi^{2}$ distribution with k=N1k=N-1 degrees of freedom if all the parameters are known, and its density function is given by,
@@
$F_{\chi^{2}}(z)=\int_{0}^{z}f_{\chi^{2}}(x)dx.$ (16)
Figure 3: Probability density function (15) for $\chi^{2}$ distribution generated using Python packages scipy and matplotlib. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by light shed (pink color); and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dark as well as light sheds (blue as well as pink colors). Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
Figure 3 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by light shed (pink color), and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dark as well as light sheds (blue as well as pink colors). The area of the critical region is given by
@@
If the typical values of the parameters $\mu,T$ in Eq. (47) are known, we have k=N1k=N-1. However, if they are unknown, in that case, k=N2k=N-2. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if the $p$-value is less than $\alpha$.
Thus, our two-fold statistical method ensures the distribution of the photon statistics to be sub-Poissonian, Poissonian, or super-Poissonian up to a certain confidence level. As in most cases, the equality of the mean and the variance denotes a Poisson distribution; for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test to ensure the completeness of the discussion.
@@
Methods Distributions with mean $m$ = variance $\sigma^{2}$
Geometric Poisson Normal
Dataset 0 Dataset 1 Dataset 2 Dataset 3
($p=0.5$) ($m=0.5$) ($m=10$) ($m=\sigma^{2}=0.5$)
Direct comparison 0.5% 4.2% 0.0% 3.3%
Comparison and GoF 100% 100% 100% 100%
Dispersion 81.3% 98.1% 98.2% 90.8%
Dispersion and GoF 100% 99.4% 99.5% 100%
Our Method (Phase-I) 85.3% 98.9% 99.1% 94.0%
Our Method (Phase-II) 100% 99.4% 99.4% 100%
Table 1: Comparison for the success rates of different methods to identify the nature of the underlying distributions, all having equal mean and variance. The methods without a goodness-of-fit (GoF) test are capable only of concluding whether the mean and the variance are sufficiently close. Given that the mean and the variance are found to be close enough, the GoF test (Phase-II) confirms whether the underlying distribution is Poisson or not. The dispersion method Fisher1992 ; Alves2022 based on the dispersion index identifies the distribution with high accuracy. However, our proposed method outperforms these methods. Note that a success rate of 0% indicates that the corresponding method fails to correctly predict the underlying distribution for all samples in the particular dataset. Similarly, a success rate of 100% indicates the correct prediction by the corresponding method for all samples in the particular dataset.
III.4 Experimental Results
We have experimented to validate our method with synthetic datasets generated by Python 3 10.5555/1593511 . We have generated one dataset from geometric distribution (3) with $p=0.5$, and one dataset from normal distribution with $m=\sigma^{2}=0.5$. Also, we have considered two different means, 0.5 and 10, to generate two sets of data for the Poisson distribution. We have used the numpy library harris2020NumPy to generate 1000 samples, each of size 100K, for each of the datasets. Table 1 shows the performance comparison between the existing methods and our proposed method with $99\%$ confidence. The methods without a goodness-of-fit (GoF) test are capable only of concluding whether the mean and the variance are sufficiently close. If they are found close in $s$ out of $S$ (= 1000 in our case) samples, the success rate of Phase-I is given by $\frac{s}{S}$. The GoF test (Phase-II) is run on these $s$ samples to confirm whether the underlying distribution is Poisson or not. If it correctly predicts Poisson or non-Poisson in $t$ out of the above $s$ samples, the success rate of Phase-II is given by $\frac{t}{s}$.
For example, as shown in Table 1, the direct comparison method for Dataset 0, generated from a geometric distribution, yields $0.5\%$ success rate. It means that in only $0.5\%$ of the 1000 samples, the direct comparison method concludes that the mean and the variance are close enough to decide that they are equal with $99\%$ confidence. Among these $1000\times 0.5\%=5$ particular samples, the goodness-of-fit test for Poisson decides for $100\%$ of them (i.e., for all the above 5 samples) that the data is not generated from a Poisson distribution.
On the other hand, for Dataset 2 generated from the Poisson distribution, our interval estimation process (Phase-I) decides that the mean and the variance are close enough in $99.1\%$ of the samples with $99\%$ confidence. Among these $1000\times 99.1\%=991$ particular samples, the goodness-of-fit test for Poisson decides for $99.4\%$ of them (i.e., for 985 of the above 991 samples) that the data is indeed generated from a Poisson distribution.